RC Meeting, Wednesday, October 14 , 2021
RC Meeting Information
RC Daily Meetings are at 10 am, remote connection link: https://bluejeans.com/861439920
- Very efficient beam delivery over the last 24 hours
- Hall A pass change shown as taking place on Monday on the white board --> this may change depending on commissioning progress. Hall A will know more once SBS comes up.
- Next week there will be a maintenance day (Wednesday on white board). May be the possibility for access to the hall.
Hall C Status
- Efficient data taking - no significant hall-related downtime
- So far we have collected 1290 pions (as of 7:40 am) in this configuration
- One large excursion in bender LHe level . These excursions seem to happen every couple days 
Current Run Plan
- Continue data taking until early Nov
- Replace flow switches for magnets: next opportunistic access (Joe Beaufait) ( PENDING )
- The HB has seen some fluctuations in the LHe levels. At the next opportunistic access Heidi would like to investigate the cryo line. ( PENDING )
- Heidi believes this is related to a vacuum issue and the access is for diagnostic purposes only. Can be escorted during regular office hours (do not call her in at midnight for this)
- If indeed it proves to be the case the only remedy would be to warm up the magnet and pump the region (which would be a multiple day operation)
- So far the PID loops have done their job and Heidi (together with the shift crew) are keeping a close eye on the LHe levels
- Water flow to the dump aperture needs to be replaced: needs 4+h and Steve Suhring believes that it can wait until the SAD ( PENDING )
- Target ladder camera was identified to be bad: Joe Beaufait will purchase a replacement in the next funding cycle ( PENDING )
- Still need to follow-up with Joe about when the purchase can be done.
- Still running with no UPS on the SHMS cryo controls. ( PENDING )